(1 point) Find the length of the curver r(t) = i +3tj + tºk, 0<t</96 L
Answer #1

The The given curve olt) = i tttt oct=j96 REBNO are length 06 given given interval (a,b] is given by La poticos de curve in qkhen to then u= 0+4=4 then u= (146)²+4 =100 t= 396 tou of ra du = 3/243727100 is 3 J4 4 = = [us] yous 4 [ Clouse - (939) [ 10